OpenStudy (yanasidlinskiy):

How would I solve this type of equation? Find the function f(x) - x + 9, find (f o f-1)(5) A. 14 B. 5 C. 25 D. -5

OpenStudy (mathmale):

Hello, Yana! Here you have the composition of function f(x) with itself. This means that f(x) is the input to f(x): f(f(x)). Do that algebraically first. When done, let x=5.

OpenStudy (yanasidlinskiy):

Ok. So would it be: f(x)-5+9?

OpenStudy (mathmale):

Yana, you haven't yet taken the composition f(f)x)), so, no. Try again. If you define f(x) as f(x)=-x+9 (check that, please), then f( f(x) ) = f(-x+9) = -(-x+9)+9=??

OpenStudy (mathmale):

just simplify the algebra here.

OpenStudy (yanasidlinskiy):

Ok. You want me to show you what I got?

OpenStudy (mathmale):

You bet I do!

OpenStudy (yanasidlinskiy):

Ok. I'm a little confused. What do you mean by taking out the composition?

OpenStudy (anonymous):

Still facing problem??

OpenStudy (yanasidlinskiy):

Yes. I do. Sorry couldn't reply earlier.

OpenStudy (yanasidlinskiy):

The problem is, I don't know how to set up the equation.

OpenStudy (anonymous):

Where exactly??

OpenStudy (anonymous):

Okay, it is okay, you should not be worried instead you must be happy that today you are going to add one more new thing to your knowledge.. :)

OpenStudy (yanasidlinskiy):

This part: (f o f-1)(5) <-----Is that where I'm supposed to plug in, f(x) - x + 9?? That's where I'm confused.

OpenStudy (yanasidlinskiy):

Haha, I hope I understand this:D

OpenStudy (anonymous):

Is that f^{-1} there??

OpenStudy (yanasidlinskiy):

Yes.

OpenStudy (anonymous):

Firstly, I want to tell you about function somewhat..

OpenStudy (aum):

One point that may help solve this problem quickly is: If f(x) is a function and g(x) is its inverse then, f(g(x)) = x and g(f(x)) = x

OpenStudy (anonymous):

Suppose, you have f(x) = x + 3, here f(x) means a function of x, as you can see here, f(x) value depends on x (3 is constant there)..

OpenStudy (anonymous):

I mean to say that: f(x) = x + 3 If I put x = 1,2,3 and so on, f(x) value will also change accordingly.. Here in f(x), x is basically called argument... At the starting, you must have confusion, but as we proceed, you can easily get what we are doing..

OpenStudy (yanasidlinskiy):

Ok, yea. It's starting to make sense:D

OpenStudy (anonymous):

One more thing in advance I want to tell you that : f(x) = x+3 Now if I change my argument from x to (x+9) then where are you seeing x earlier, just replace it with (x+9) So, I am replacing all x with (x+9) Look it here, we get as: f(x+9) = (x+9) + 3 Look it very carefully, are you getting it?? Look 100 times and say are you getting it because this is the point where you will be getting confused I suppose..

OpenStudy (yanasidlinskiy):

Yes!!!!!!!!!!!!!!!!!!!!!!!!! That's the part!!!!!!!!!!!!!!!!!!!!! Wow!!!! Youre such a great explainer!! By the way..are you a teacher? Or somethin like that?

OpenStudy (anonymous):

Are you a student??

OpenStudy (yanasidlinskiy):

Yes.

OpenStudy (anonymous):

I will turn 90 this 5th Of August.. :)

OpenStudy (yanasidlinskiy):

Wow!!! Ok. Anyways. Back on track:) And Happy Early Birthday:D

OpenStudy (anonymous):

I think you will oppose that, but anyway, back on track now and don't distract me.. :P

OpenStudy (yanasidlinskiy):

Haha:) Ok. I won't hopefully:D

OpenStudy (anonymous):

So getting that part??

OpenStudy (anonymous):

Now, some more you have to learn before proceeding to our question.. f(x) can be written as y.. Okay?? So, f(x) = x+3 or y = x+3 both are one and the same thing..

OpenStudy (yanasidlinskiy):

Ok. Yes. I'm understanding it soo far.:D

OpenStudy (anonymous):

Now, I want to know if you know how to know I mean how to find Inverse of f(x) ??

OpenStudy (yanasidlinskiy):

Yes. I think so.

OpenStudy (anonymous):

Should I give an example or is it okay??

OpenStudy (yanasidlinskiy):

Give an example. I think that would be better:D

OpenStudy (anonymous):

So, let us take a simple function here to understand how to find inverse.. Let us take f(x) = x+5 As I said, you can write f(x) as y too, so : y = x+ 5 Now First step would be to convert this expression in terms of x.. So, subtracting 5 from both sides, y - 5 = x, (this is in now terms of x) Now second step is to just interchange x with y and y with x.. So, y = x - 5, and this y is here \(f^{-1}(x) = x-5\) Check it till here if I am not wrong then..

OpenStudy (anonymous):

Is it okay or somewhat hard to understand??

OpenStudy (yanasidlinskiy):

Umm..Will I be able to do it similarily with this one too? f(x) - x + 9,

OpenStudy (anonymous):

Yeah why not...

OpenStudy (anonymous):

So, let us do first this only, okay??

OpenStudy (yanasidlinskiy):

Ok:) Can I show you what I have done by the way you have it and tell me if it's right or wrong?

OpenStudy (anonymous):

Yeah, that will be better.. I am spectator now, Go Ahead.. :P

OpenStudy (yanasidlinskiy):

Hahha:D Ok. Umm. I'm not sure if I got this right, but I have: \[f ^{-1}(x)=x-9\]

OpenStudy (anonymous):

Wait, wait sorry to interrupt in between, In question: f(x) = x+9 no??

OpenStudy (anonymous):

I mean in between that's equal to sign no??

OpenStudy (anonymous):

What do you think, have you found correctly for inverse of f(x) ??? YES or NO??

OpenStudy (yanasidlinskiy):

There is no equal sign.

OpenStudy (anonymous):

you mean f(x) - x + 9 ???

OpenStudy (yanasidlinskiy):

Yes. That's what I have above in my question.

OpenStudy (anonymous):

Yana, this will create a problem, just look this problem once again, for sure f(x) = x+9 must be correct..

OpenStudy (yanasidlinskiy):

I have it both ways. The one that we're talking about and f(x)=x+9. Lets just do \(\huge{\leftarrow}\)

OpenStudy (anonymous):

I did not get you??

OpenStudy (anonymous):

So, back to question, And yes the inverse you have found is correct (as I am the teacher here, just kidding), Good, you are learning something.. :)

OpenStudy (anonymous):

Hey, I guess you are not writing a Novel??

OpenStudy (yanasidlinskiy):

I have to solve \(\huge{f(x) - x + 9}\) and \(\huge{f(x)=x+9}\) They are totally two different questions. For the second one I have: \(\huge{f(x)=x-9}\)

OpenStudy (yanasidlinskiy):

Hhaa. No. I wasn't.

OpenStudy (anonymous):

What does - sign mean there??

OpenStudy (yanasidlinskiy):

Negative.

OpenStudy (anonymous):

Then I think equal to sign is missing there.. You have two question as : f(x) = -x + 9 and f(x) = x+9, these two will make sense..

OpenStudy (yanasidlinskiy):

Lets just stick with: \(\huge{f(x)=x+9}\)

OpenStudy (anonymous):

Right..

OpenStudy (mathmale):

Yana: Apologies for not having stuck around to help you finish this problem. I can't help but wonder if the f(x) - x + 9 you've typed in should be f(x) = -x + 9. Would you please check that?

OpenStudy (anonymous):

As you have found its inverse, so now we should enter the territory of desired result..

OpenStudy (anonymous):

@mathmale you can start it from here if you want, I have no problem... :)

OpenStudy (mathmale):

To get anywhere with this problem, we MUST have f(x) definied somewhere. I don't see that. f(x) - x + 9 is not a definition of a function; f(x)=-x+9 is.

OpenStudy (mathmale):

@waterineyes: Thank you so much for all the effort you've put into this.

OpenStudy (anonymous):

Any OS user will do the same.. :)

OpenStudy (yanasidlinskiy):

I promise you. The question doesn't have an equal sign.

OpenStudy (anonymous):

Yana, there can be omission too, printing mistake can be there..

OpenStudy (anonymous):

A function is defined by f(x) = x terms like this, It is its format...

OpenStudy (anonymous):

Now let us stick to f(x) = x+9, this we will do for you, and f(x) = -x + 9 this you will do for yourself only.. :P

OpenStudy (yanasidlinskiy):

Lets just stick with this equation first like I have said above: \(\huge{f(x)=x+9}\) And my answer is: \[f ^{-1}(x)=x-9\]

OpenStudy (yanasidlinskiy):

It's Ok @mathmale Not everyone can do it.

OpenStudy (anonymous):

Now hurry up.. So, we have with ourselves : \(f^{-1}(x) = x - 9 \)

OpenStudy (yanasidlinskiy):

Would I plug in 5 for x?

OpenStudy (anonymous):

Wait, you have f(x) and f^{-1}(x), in which x you are trying to plug in??

OpenStudy (yanasidlinskiy):

x-9

OpenStudy (anonymous):

If you want to learn more then: \[(f \circ f)x \implies f[f(x)]\]

OpenStudy (anonymous):

Similarly: \[(f \circ f^{-1}) \implies f[f^{-1}(x)]\]

OpenStudy (anonymous):

So, here you are given with: \[(f \circ f^{-1})(5) \implies \color{green}{f[f^{-1}(5)]}\]

OpenStudy (yanasidlinskiy):

Ok. So, what would I do next?

OpenStudy (anonymous):

You know : f^{-1}(x) = x - 9 and to find f^{-1}(5) you will replace x with 5, so go ahead..

OpenStudy (anonymous):

Getting or not??

OpenStudy (yanasidlinskiy):

Yes. Let me type out what I have.

OpenStudy (anonymous):

You can write it without LATEX too because I don't want you to start writing Novel again.. :P

OpenStudy (yanasidlinskiy):

Hahhaa!!:D sorry if it really does take that long. \[f ^{-1}(x)=5-9\]

OpenStudy (anonymous):

You can subtract it too otherwise you should have calculator with you if you don't want your brain do this calculation.. :P

OpenStudy (anonymous):

And please note since you are replacing x with 5 so replace it on both LHS and RHS.. \(f^{-1}(x) = x - 9 \implies f^{-1}(5) = 5- 9\) like this..

OpenStudy (yanasidlinskiy):

I would plug that into my calculator. Right?

OpenStudy (anonymous):

Are you serious??

OpenStudy (anonymous):

5 - 9 = ??

OpenStudy (yanasidlinskiy):

Hahha. I was just kidding. I know what is is. It's -4:D

OpenStudy (anonymous):

Don't do that otherwise I will tell moderators to investigate how you got to 75.. :P

OpenStudy (yanasidlinskiy):

Omg. Are you for real? What about that \(\huge{\uparrow}\) 90 years old?

OpenStudy (mathmale):

\[(f \circ f^{-1})(5) \implies \color{green}{f[f^{-1}(5)]}\] I'd have to check that out; not sure that I'd agree. I'd much prefer you write \[f(f ^{-1}(x))\]FIRST and substitute x=5 LAST. No, Yana, you do not need a calculator here. If you start with ANY f)x) and find the inverse of f(x), then the composition of the two separate functions MUST equal x. In fact this is the test for correctness when you are finding the inverse of a function. Thus, if you are really ending up with f(x) and \[f ^{-1}(x), \] the composition must be and is simply x. So there we let x=5 and we're done.

OpenStudy (anonymous):

So our expression further reduces to: \[f[f^{-1}(5)] \implies \color{red}{ f[-4]}\]